OpenStudy (anonymous):

Linear Equations - Solve each equation or formula for the variable indicated. - 7h + f = 2h + g for g.

OpenStudy (jackellyn):

Your goal is to get the g alone. How can you do that?

OpenStudy (anonymous):

-2h from -7h on both sides?

OpenStudy (jackellyn):

You're right about -2h on both sides but notice that the 7h is positive 7h-2h+f=g and now you combine like terms, so what is the answer @Absullum457 ?

OpenStudy (anonymous):

5h = G/F?

OpenStudy (jackellyn):

close, the g has to be alone. so leave the f where it is

OpenStudy (anonymous):

5h/F = G?

OpenStudy (anonymous):

Or is the equation already solved?

OpenStudy (jackellyn):

Its already solved, you dont have to do anything to the f. So you get 5h+f=g

OpenStudy (anonymous):

OHHH okay I get it now. Much thanks to you.

OpenStudy (jackellyn):

No problem :) Good work!